Hey, I'd be happy to get a solution for this exercise:
The product of four prime numbers (they don't necessarily need to be all different) is the same as ten times the sum of these prime numbers.
Define all possibiliies for these four prime numbers.
THANK YOU REALLY MUCH

Answer by Edwin McCravy(20059)


2×3×5×5 = 150 and also 10(2+3+5+5) = 10(15) = 150
I don't think there are any more cases.
What do you mean, "define all possibilities"??
Edwin